23,364 in 1990
23,226 in 1995
25,088 in 2000
What is:
a. the percent increase from 1990 to 1995.
b. the percent increase from 1995 to 2000.
c. the percent increase from 1990 to 2000.
please explain how you get the answer is very simple terms, thank you.


Answer by mukhopadhyay(490)

Percent increase from 1990 to 1995
= 100[(23226-23364)/23364] = -13800/23364 = -.6% (approx)... it is a decrease
.........
Percent increase from 1995 to 2000
= 100[(25088-23226)/23226] = 186200/23226 = 8% (approx)
..........
Percent increase from 1990 to 2000
= 100[(25088-23364)/23364] = 172400/23364 = 7.4% (approx)